The waste management industry reports that Canadian individuals dispose of an average of 760 kilograms of nonhazardous material in a given year. Assume a standard deviation of 90 kilograms. Find the probability that the mean of a sample of 55 Canadians will be between 750 and 774 kilograms.
Answer by Boreal(15235)

t df=55=(x bar-mean)/s/sqrt(n)
=(774-760)*sqrt(55)/90=+1.15, top value
=(750-760)*sqrt(55)/90=-0.82, lower value
0.6646 is the probability.
If it were a z-test (check), the std error of the mean would be 12.1, and this would be very close to +/- 1 sd from the mean or 68%.